OpenStudy (anonymous):

Solve equation: x-2y=3 5x-y=6

hartnn (hartnn):

multiply equation 1 by 5 and subtract the two equations....this should help...

OpenStudy (waleed_imtiaz):

multiply the first equation by 5... and then subtract it from the second..... U will get 5x-10y-15-5x+y+6=-9y=-9.... y=1..put y=1 in any of the equation u will get x...........Solved

OpenStudy (anonymous):

if i'm correct than x=5

hartnn (hartnn):

can u show your work,because its incorrect :(

OpenStudy (waleed_imtiaz):

y=-1..now solve for x

OpenStudy (anonymous):

x-2y=3 x-2(1)=3 x-2=3 x=2+3 x=5

hartnn (hartnn):

but u got y=-1??

OpenStudy (anonymous):

yep, i messed that up...x=1

hartnn (hartnn):

nice :) u got it now...

OpenStudy (anonymous):

thank you both